# Beyond Instruction Following: Evaluating Inferential Rule Following of Large Language Models

### v2

@@ Abstract
Although Large Language Models (LLMs) have demonstrated strong instruction-following ability, they are further supposed to be controlled and guided by rules in real-world scenarios to be safe, accurate, and intelligent. This demands the possession of inferential rule-following capability of LLMs. However, few works have made a clear evaluation of the inferential rule-following capability of LLMs. Previous studies that try to evaluate the inferential rule-following capability of LLMs fail to distinguish the inferential rule-following scenarios from the instruction-following scenarios. Therefore, this paper first clarifies the concept of inferential rule-following and proposes a comprehensive benchmark, RuleBench, to evaluate a diversified range of inferential rule-following abilities. Our experimental results on a variety of LLMs show that they are still limited in following rules. Our analysis based on the evaluation results provides insights into the improvements for LLMs toward a better inferential rule-following intelligent agent. We further propose Inferential Rule-Following Tuning (IRFT), which outperforms IFT in helping LLMs solve RuleBench. The data and code can be found at: https://anonymous.4open.science/r/llm-rule-following-B3E3/
[…]
@@
Wangtao Sun1,2, Chenxiang Zhang1, Xueyou Zhang1, Ziyang Huang1,2, Haotian Xu6, Pei Chen4, Shizhu He1,2, Jun Zhao1,2, Kang Liu1,2,5Corresponding author: Kang Liu (kliu@nlpr.ia.ac.cn) 1The Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences, Beijing, China 2School of Artificial Intelligence, University of Chinese Academy of Sciences, Beijing, China 3CAS Engineering Laboratory for Intelligent Industrial Vision, Institute of Automation, Chinese Academy of Sciences, Beijing, China 4Department of Computer Science and Engineering, Texas A&M University 5Shanghai Artificial Intelligence Laboratory 6Xiaohongshu Inc sunwangtao2021@ia.ac.cn
@@ 1 Introduction
Benefiting from a vast amount of pre-training data and the enormous parameters, the Large Language Models (LLMs) can accomplish numerous Natural Language Processing (NLP) tasks thanks to their instruction-following ability. However, in real-world applications, people often expect LLMs to generate outputs that conform to various rules. Humans use rules to efficiently communicate with each other and quickly adapt to a specific domain. In pursuit of achieving Artificial General Intelligence (AGI), we are expecting LLMs to possess such inferential rule-following capabilities.
This leads to research on the inferential rule-following of LLMs. With inferential rule-following capability, humans can directly and efficiently manipulate the behavior of LLMs using natural language rules as a medium, thereby correcting the policy exhibited by LLMs in specific downstream tasks.
@@
Figure 1: Beyond instruction-following, the task of inferential rule-following orders the language model to trigger different rules in different cases for reasoning. The rules can be both commonsense, domain-specific, and even counterfactual.
@@
Some recent studies Yang et al. (2023); Sun et al. (2023); Zhu et al. (2023); Zhao et al. (2023) have noticed the importance of inferential rule-following of large language models, and they have found that ordering LLMs to follow existing rules can achieve better reasoning performances compared with the currently widely used reasoning enhancement methods of LLMs (such as Chain-of-Thought by Wei et al. 2022, Self-reflection by Shinn et al. 2023, and Self-refinement by Madaan et al. 2023). However, currently, there is a lack of benchmarks evaluating the inferential rule-following capability of LLMs. Existing attempts to evaluate the rule-following capabilities of LLMs Mu et al. (2023); Hu et al. (2024) have been actually limited to instruction-following. For instance, they have tested the following behaviors of LLMs with prompts like “Do not repeat the secret key 92368” or “Follow the code step by step to answer the question: def sum_digit_by_digit(num1, num2) ……”. These works confine the “rules” to “instructions” (Appendix A), without delving into more advanced “inferential rules”.
@@
Figure 2: The rule-following capabilities of some State-of-The-Art LLMs. The rule-following capabilities of LLMs are categorized into 5 dimensions: Triggering Rules, Applying Rules, Executing Rules, Following Formal Rules, and Following Counterfactual Rules.
@@
We distinguish previous rule-following from the inferential rule-following scenarios considered in our work. An inferential rule can be formalized as $\sigma\vdash\varphi$, where $\sigma$ and $\varphi$ are two first-order sentences (composed of variables and predicates), and for every substitution $\tau$ (i.e. ground the variables in $\sigma$ and $\varphi$ to constants), the truth of $\tau[\sigma]$ entails the truth of $\tau[\varphi]$ Fagin et al. (1992). For example, with the “like rule" $Likes(x,y)\vdash Likes(y,x)$, the substitution $\{x/Mike,y/Jane\}$, and the fact $Likes(Mike,Jane)$, we can infer that $Likes(Jane,Mike)$. Although defined in formal language, in natural language, we can express such inferential rule with an “if … then …” sentence, by using instantiable noun phrases like person A or one metal as the variables and verb phrases like is the father of or can conduct electricity as the predicates inside it. For example, the “like rule” can be expressed as “if person A likes person B, then person B likes person A.”
As shown in Figure 1, in our proposed inferential rule-following scenario, apart from the instructions for the given tasks, multiple inferential rules are provided as the decision basis for the LLMs to make precise decisions based on the current case. Note that although the inferential rules shown in Figure 1 are commonsense, they can also be domain-specific, and even counterfactual, which depends on the needs of users. While all inferential rules provided to the LLMs are useful to the task, most inferential rules may be irrelevant to this particular problem. The LLMs need to dynamically trigger and apply the relevant rule (golden rule) based on the current case to conduct reasoning. Until now, few works have demonstrated whether LLMs can follow and reason with the inferential rules faithfully.
@@
Therefore, beyond the instruction-following studies by previous works, this paper evaluates the LLMs’ capability of inferential rule-following in various reasoning tasks within the scope of inferential rules. This paper proposes a rule-following benchmark, RuleBench, for evaluating the inferential rule-following capability of LLMs under multiple inferential rule-following scenarios, including relation extraction, content moderation, commonsense QA, science QA, and judgment prediction. Based on RuleBench, this paper has evaluated multiple State-of-The-Art LLMs (§4.1), discusses the impact of different rule quantities (§4.2), rule forms (§4.3), Chain-of-Thought (CoT) in applying rules (§4.4), and explores whether the capability of inferential rule-following still exists in counterfactual scenarios (§4.5). This paper also analyzes the cases where LLMs fail to follow the rules (§4.6), categorizing them into Triggering Error and Execution error, which stand for the cases where LLMs fail to trigger the golden rule and LLMs fail to execute the golden rule, respectively. Based on the results, as shown in Figure 2, we categorize the inferential rule-following capabilities of LLMs into 5 dimensions (§4.7), including the capabilities of Triggering Rules, Applying Rules, Executing Rules, Following Formal Rules, and Following Counterfactual Rules. Finally, to further improve the inferential rule-following capabilities of LLMs, we propose the Inferential Rule-Following Tuning (IRFT) that enables LLMs to learn to trigger and apply the correct inferential rule based on the current cases (§4.8). The experimental results show that IRFT significantly outperforms IFT. In summary, the major contributions of this paper are as follows:
@@
  • We evaluated the capabilities of inferential rule-following of multiple State-of-The-Art LLMs on various tasks and rule settings, and categorized their inferential rule-following abilities into 5 dimensions. Based on the results, we analyze the possible reasons that limit the inferential rule-following capabilities of current LLMs and provide some insights into the improvements for LLMs toward a better inferential rule-following intelligent agent.
  • We propose the Inferential Rule-Following Tuning (IRFT) that enables LLMs to learn to trigger and apply the correct inferential rule based on the current cases. The experimental results show that IRFT significantly outperforms IFT.
@@ 2.1 Rule-enhanced LLM Reasoning
While LLMs have demonstrated remarkable zero-shot reasoning capabilities in many downstream tasks, they still generate outputs that do not conform to logic or human preference. Some research studies have found that compared with the reasoning enhancement methods based on LLMs themselves like Chain-of-Thought Wei et al. (2022), Self-reflection Shinn et al. (2023), and Self-refinement Madaan et al. (2023), providing LLMs with relevant rules with Retrieval-Augmented Generation (RAG) paradigm do better in helping them conduct reasoning in the downstream tasks Yang et al. (2023); Sun et al. (2023); Zhu et al. (2023); Zhao et al. (2023). However, the inferential rule-following capability of LLMs is far from satisfactory. Few works have comprehensively evaluated whether LLMs can benefit from the provided rules under different scenarios and how LLMs can follow the rules better. To make up for this gap, this paper conducted a series of experiments to evaluate the inferential rule-following capabilities of several State-of-The-Art LLMs and provide some insights into how LLMs can follow rules better.
@@ 2.2 LLMs Instruction-following
Instruction-following has been generally considered an important capability of LLMs Zhong et al. (2021); Mishra et al. (2021); Wei et al. (2021); Yin et al. (2023), and some previous works have been done to evaluate the instruction-following capability of LLMs Zhou et al. (2023); Qin et al. (2024). However, only a few works have cast their attention to the question of inferential rule-following. Recent works focused on the rule-following capability of LLMs Mu et al. (2023); Hu et al. (2024) confined the rule-following to instruction-following. Instead, this paper proposes the scenario of inferential rule-following and sets up useful baselines for future works.
@@ 3 RuleBench
Under the scenarios introduced above, As shown in Figure 3, RuleBench involves multiple settings of inferential rule-following, to comprehensively evaluate the LLMs from different perspectives. The settings include rule quantity (i.e. how many rules are provided to the LLMs while only one of them is relevant to the current case), rule form (i.e. which form the rules illustrated in, natural language or formal language), the presence of Chain-of-Thought when applying rules (i.e. directly generate the answer based on the question and rules, or trying verbally apply the rule to the question before answering it), and rule factuality (i.e. whether the conclusion of the rule is factual or counterfactual). RuleBench allows us to analyze the failure cases of inferential rule-following from a behavioral perspective, classifying them into Triggering Error (i.e. LLMs fail to trigger the golden rule) and Execution Error (i.e. LLMs success to trigger the golden rule but fail to execute the golden rule).
@@ 4 Evaluation
To comprehensively evaluate the inferential rule-following capabilities of LLMs, based on the proposed RuleBench, this paper has designed 5 main parts of experiments. We evaluate the effects of rule quantity (§4.2), rule form (§4.3), the presence of CoT when applying rules (§4.4), and rule factuality (§4.5). Besides, we analyzed the failure cases of inferential rule-following from a behavioral perspective, classifying them into Triggering Error and Execution Error4.6). Based on these evaluation results, we categorize the inferential rule-following capabilities into 5 dimensions and compare the performances of 8 State-of-The-Art LLMs (§4.7). Finally, to further improve the inferential rule-following capabilities of LLMs, we propose the Inferential Rule-Following Tuning (IRFT) that enables LLMs to learn to trigger and apply the correct inferential rule based on the current cases (§4.8).
@@ 4.1 Model Selections
For open-source LLMs, we adopt Llama-2-7b-chat Touvron et al. (2023), Meta-Llama-3-8B AI@Meta (2024), Mistral-7B-Instruct-v0.2 Jiang et al. (2023), YiYoung et al. (2024), and Phi-3Abdin et al. (2024). For closed-source LLMs, we adopt gpt-3.5-turbo, gpt-4-turbo OpenAI (2023), and gpt-4o from OpenAI. The comprehensive performance comparison of them is shown in Figure 2 and the explanation and analysis is in §4.7.
@@
Figure 5: The inferential rule-following performance of LLMs with rules of formal language (FOL) and natural language (NL).
Figure 6: The inferential rule-following performance of LLMs when applying rules with or without using Chain-of-Thought.

### v4

@@ Abstract
Although Large Language Models (LLMs) have demonstrated strong instruction-following ability, they are further supposed to be controlled and guided by inferential rules in real-world scenarios to be safe, accurate, and intelligent. This demands the possession of inferential rule-following capability of LLMs. However, no prior work has made a clear evaluation of the inferential rule-following capability of LLMs. Previous studies that try to evaluate the inferential rule-following capability of LLMs fail to distinguish the inferential rule-following scenarios from the instruction-following scenarios. Therefore, this paper first clarifies the concept of inferential rule-following and proposes a comprehensive benchmark, RuleBench, to evaluate a diversified range of inferential rule-following abilities. Our experimental results on a variety of LLMs show that they are still limited in following rules. Our analysis based on the evaluation results provides insights into the improvements for LLMs toward a better inferential rule-following intelligent agent. We further propose Inferential Rule-Following Tuning (IRFT). The experimental results show that through IRFT, LLMs can learn abstract inferential rule-following abilities from purely synthetic data and then generalize to RuleBench. The data and code can be found at: https://anonymous.4open.science/r/llm-rule-following-B3E3/
@@
Wangtao Sun1,2, Chenxiang Zhang1, Xueyou Zhang1, Xuanqing Yu3,7, Ziyang Huang1,2, Haotian Xu6, Pei Chen4, Shizhu He1,2, Jun Zhao1,2, Kang Liu1,2,5Corresponding author: Kang Liu (kliu@nlpr.ia.ac.cn) 1The Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences, Beijing, China 2School of Artificial Intelligence, University of Chinese Academy of Sciences, Beijing, China 3CAS Engineering Laboratory for Intelligent Industrial Vision, Institute of Automation, Chinese Academy of Sciences, Beijing, China 4Department of Computer Science and Engineering, Texas A&M University 5Shanghai Artificial Intelligence Laboratory 6Xiaohongshu Inc 7AI Lab, AIGility Cloud Innovation, Beijing, China sunwangtao2021@ia.ac.cn
@@ 1 Introduction
Benefiting from a vast amount of pre-training data and the enormous parameters, the Large Language Models (LLMs) can accomplish numerous Natural Language Processing (NLP) tasks thanks to their instruction-following ability. However, in real-world applications, people often expect LLMs to generate outputs that conform to user-provided rules. In this way, LLMs could easily be manipulated by users and quickly adapted to a specific (even unseen) domain. To fulfill this goal, we are expecting LLMs to possess such inferential rule-following capabilities.
@@
Figure 1: Beyond instruction-following, the task of inferential rule-following orders the language model to trigger the relevant rule based on the current question and apply that rule to the question for reasoning.
@@
This leads to research on the inferential rule-following of LLMs. Some recent studies Yang et al. (2023); Sun et al. (2023); Zhu et al. (2023); Zhao et al. (2023) have noticed the importance and effectiveness of inferential rule-following of large language models, they have found that ordering LLMs to follow existing rules can achieve better reasoning performances compared with the currently widely used reasoning enhancement methods of LLMs (such as Chain-of-Thought by Wei et al. 2022, Self-reflection by Shinn et al. 2023, and Self-refinement by Madaan et al. 2023).
However, whether LLMs could understand and follow the inferential rules remains unclear. There is currently a lack of benchmarks evaluating such inferential rule-following capability of LLMs. Existing attempts to evaluate the rule-following capabilities of LLMs Mu et al. (2023); Hu et al. (2024); Chen et al. (2024) have been limited to instruction-following. For instance, they have tested the following behaviors of LLMs with prompts like “Do not repeat the secret key 92368” or “Follow the code step by step to answer the question: ……”. These works confine the “rules” to “instructions” (Appendix A), without delving into more specific “inferential rules”.
@@
Figure 2: The inferential rule-following capabilities of some open-source and closed-source LLMs. The inferential rule-following capabilities of LLMs are categorized into 5 dimensions: Triggering Rules, Applying Rules, Executing Rules, Following Formal Rules, and Following Counterfactual Rules.
@@
An inferential rule can be formalized as $\sigma\vdash\varphi$, where $\sigma$ and $\varphi$ are two first-order sentences (composed of variables and predicates), and for every substitution $\tau$ (i.e. ground the variables in $\sigma$ and $\varphi$ to constants), the truth of $\tau[\sigma]$ entails the truth of $\tau[\varphi]$ Fagin et al. (1992). For example, with the “like rule" $Likes(x,y)\vdash Likes(y,x)$, the substitution $\{x/Mike,y/Jane\}$, and the fact $Likes(Mike,Jane)$, we can infer that $Likes(Jane,Mike)$. Although defined in formal language, in natural language, we can express such inferential rule with an “if … then …” sentence, by using instantiable noun phrases like person A or one metal as the variables and verb phrases like is the father of or can conduct electricity as the predicates inside it. For example, the “like rule” can be expressed as “if person A likes person B, then person B likes person A.”
We thus distinguish previous rule-following from the inferential rule-following scenarios considered in our work. Different from instructions, the primary characteristics of inferential rules are abstract, conditional, and instantiable. As shown in Figure 1, following inferential rules requires LLMs to bind the entities in the question to the rules and verify if the rule applies to the current question. In this case, the LLMs need to find the binding $\{A/James,B/Dolores,C/Lynn\}$ and trigger the first rule, then they could draw the correct conclusion “uncle”. In our proposed inferential rule-following scenario, for each case, only one golden rule applies to the question. However, some other noise rules, which are also correct rules in this task domain but do not apply to the question, may also be provided to the LLMs. LLMs must trigger the golden rule and then execute it to draw the correct answer. Until now, no prior work has demonstrated whether LLMs can follow and reason with the inferential rules faithfully.
@@
Therefore, beyond the instruction-following studies by previous works, this paper evaluates the LLMs’ capability of inferential rule-following in various reasoning tasks within the scope of inferential rules. This paper proposes a rule-following benchmark, RuleBench, for evaluating the inferential rule-following capability of LLMs under multiple inferential rule-following scenarios, including relation extraction, content moderation, commonsense QA, science QA, and judgment prediction.
Based on RuleBench, this paper has evaluated multiple open-source and closed-source LLMs (§4.1) to find out whether LLMs could understand and follow the inferential rules. In specific, we answer the following questions:
  • Are inferential rules helpful to the reasoning of LLMs? (§4.2)
  • Should inferential rules be presented in natural language or formal language? (§4.3)
  • Is LLMs with Chain-of-Thought (CoT) able to effectively apply the rules? (§4.4)
  • Can LLMs still follow inferential rules in counterfactual scenarios? (§4.5)
This paper also analyzes the cases where LLMs fail to follow the rules (§4.6), categorizing them into Triggering Error and Execution Error, which stand for the cases where LLMs fail to trigger the golden rule and LLMs fail to execute the golden rule, respectively.
Based on these results, as shown in Figure 2, we categorize the inferential rule-following capabilities of LLMs into 5 dimensions (§4.7) to help us intuitively grasp the inferential rule-following capabilities of these LLMs.
Finally, to further improve the inferential rule-following capabilities of LLMs, we propose the Inferential Rule-Following Tuning (IRFT) that enables LLMs to learn to trigger and apply the correct inferential rule based on the current cases (§5). The experimental results show that through IRFT, LLMs can learn abstract inferential rule-following abilities from purely synthetic data and then generalize to RuleBench. In summary, the major contributions of this paper are as follows:
@@
  • We evaluated the capabilities of inferential rule-following of multiple open-source and closed-source LLMs on various tasks and rule settings, and categorized their inferential rule-following abilities into 5 dimensions. Based on the results, we analyze the possible reasons that limit the inferential rule-following capabilities of current LLMs and provide some insights into the improvements for LLMs toward a better inferential rule-following intelligent agent.
  • We propose the Inferential Rule-Following Tuning (IRFT) that enables LLMs to learn to trigger and apply the correct inferential rule based on the current cases. The experimental results show that through IRFT, LLMs can learn abstract inferential rule-following abilities from purely synthetic data and then generalize to RuleBench.
@@ 2.1 Rule-enhanced LLM Reasoning
While LLMs have demonstrated remarkable zero-shot reasoning capabilities in many downstream tasks, they still generate outputs that do not conform to logic or human preference. Some research studies have found that compared with the reasoning enhancement methods based on LLMs themselves like Chain-of-Thought Wei et al. (2022), Self-reflection Shinn et al. (2023), and Self-refinement Madaan et al. (2023), providing LLMs with relevant rules with Retrieval-Augmented Generation (RAG) paradigm do better in helping them conduct reasoning in the downstream tasks Yang et al. (2023); Sun et al. (2023); Zhu et al. (2023); Zhao et al. (2023). However, the inferential rule-following capability of LLMs is far from satisfactory. No prior work has comprehensively evaluated whether LLMs can benefit from the provided rules under different scenarios and how LLMs can follow the rules better. To make up for this gap, this paper conducted a series of experiments to evaluate the inferential rule-following capabilities of several open-source and closed-source LLMs and provide some insights into how LLMs can follow rules better.
@@ 2.2 LLMs Instruction-following
Instruction-following has been generally considered an important capability of LLMs Zhong et al. (2021); Mishra et al. (2021); Wei et al. (2021); Yin et al. (2023), and some previous works have been done to evaluate the instruction-following capability of LLMs Zhou et al. (2023); Qin et al. (2024). However, only a few works have cast their attention to the question of inferential rule-following. Recent works focused on the rule-following capability of LLMs Mu et al. (2023); Hu et al. (2024); Chen et al. (2024) confined the rule-following to instruction-following. Instead, this paper proposes the scenario of inferential rule-following and sets up useful baselines for future works.
@@ 3 RuleBench
Under the scenarios introduced above, As shown in Figure 3, RuleBench involves multiple settings of inferential rule-following, to comprehensively evaluate the LLMs from different perspectives. The settings include rule quantity (i.e. how many rules are provided to the LLMs while only one of them applies to the current case), rule form (i.e. which form the rules illustrated in, natural language or formal language), the presence of Chain-of-Thought when applying rules (i.e. directly generate the answer based on the question and rules, or trying verbally apply the rule to the question before answering it), and rule factuality (i.e. whether the conclusion of the rule is factual or counterfactual). RuleBench allows us to analyze the failure cases of inferential rule-following from a behavioral perspective, classifying them into Triggering Error (i.e. LLMs fail to trigger the golden rule) and Execution Error (i.e. LLMs success to trigger the golden rule but fail to execute the golden rule).
@@ 4 Evaluation
To comprehensively evaluate the inferential rule-following capabilities of LLMs, based on the proposed RuleBench, this paper has designed 5 main parts of experiments. We evaluate the effects of rule quantity (§4.2), rule form (§4.3), the presence of CoT when applying rules (§4.4), and rule factuality (§4.5). Besides, we analyzed the failure cases of inferential rule-following from a behavioral perspective, classifying them into Triggering Error and Execution Error4.6). Based on these evaluation results, we categorize the inferential rule-following capabilities into 5 dimensions and compare the performances of 7 open-source and closed-source LLMs (§4.7). The test-time prompts used in this section can be found in Appendix C.
@@ 4.1 Model Selections
For open-source LLMs, we adopt Llama-2-7b-chat Touvron et al. (2023), Meta-Llama-3-8B AI@Meta (2024), Mistral-7B-Instruct-v0.2 Jiang et al. (2023) and Yi Young et al. (2024). For closed-source LLMs, we adopt gpt-3.5-turbo, gpt-4-turbo OpenAI (2023), and gpt-4o from OpenAI. The comprehensive performance comparison of them is shown in Figure 2 and the explanation and analysis is in §4.7.
@@
Model Form CAIL2018 SALAD TheoremQA CLUTRR ULogic DEER
All Few All Few All Few All Few All Few All Few
gpt-4o FOL 95.78 95.18 92.93 96.20 70.36 69.45 43.13 43.10 87.59 70.60 73.81 38.10
NL 96.99 97.59 95.19 95.35 69.45 69.45 58.97 62.67 94.82 79.16 92.86 71.43
Llama-2-7b-chat-hf FOL 50.60 53.01 58.27 81.22 39.16 38.18 23.00 27.48 60.36 79.04 57.14 71.43
NL 53.61 63.25 72.59 88.91 39.00 35.96 24.33 40.36 60.00 83.86 59.52 59.52
gpt-3.5-turbo FOL 84.94 86.14 73.33 96.83 49.82 53.26 41.54 53.28 63.73 66.10 71.43 78.57
NL 90.36 89.76 80.64 99.24 50.18 50.36 42.32 59.36 80.72 70.48 78.57 83.33
Yi-1.5-6B-Chat FOL 96.99 97.59 64.36 70.28 53.27 52.00 23.95 23.47 90.12 81.93 64.29 52.38
NL 96.99 95.78 75.73 80.44 52.00 49.64 17.27 21.37 87.59 82.77 73.81 66.67
gpt-4-turbo FOL 96.39 95.78 93.25 93.72 66.55 65.09 46.95 53.99 98.43 98.55 90.48 88.10
NL 96.99 98.80 94.81 94.70 66.36 64.36 49.32 53.92 98.43 98.67 88.10 90.48
Meta-Llama-3-8B FOL 89.16 93.98 77.71 77.16 48.91 50.73 38.55 41.89 92.17 97.23 83.33 80.95
NL 89.76 97.59 82.28 77.92 48.18 48.55 37.40 45.52 92.17 95.06 85.71 85.71
Mistral-7B-Instruct FOL 78.31 79.52 72.62 80.54 17.45 21.45 25.10 29.20 86.75 92.41 64.29 64.29
NL 84.34 92.77 82.36 83.43 22.18 24.55 26.62 32.73 95.42 94.46 57.14 64.29
Table 1: The of LLMs on RuleBench with rules of formal language (FOL) and natural language (NL).
Figure 5: The inferential rule-following performance of LLMs when applying rules with or without using Chain-of-Thought. The dashed box indicates the improvement (positive or negative) of w/ CoT over w/o CoT.